# Multi-modal AI for comprehensive breast cancer prognostication

### v1

@@ 2.2 The AI test is more accurate than a genomic assay in predicting cancer recurrence in HR+ patients
Figure 4: The AI test is prognostic in clinically meaningful situations. a, In triple-negative breast cancer, there is an ongoing discussion about potential strategies to de-escalate the KEYNOTE-522 regimen, for example with shorter or anthracycline-free chemotherapy regimens or complete ommitance of adjuvant treatment. While some biomarkers, such as tumor-infiltrating lymphocytes, are associated with long-term outcomes and pathological complete response, there is still a need for a more robust assessment of systemic therapy benefits. b-d, In hormone receptor-positive patients, questions about treatment selection involve the addition of adjuvant chemotherapy, extended endocrine therapy (ET), and new agents, such as CDK4/6 inhibitors. We hypothesize that high-risk HR+ patients who received adjuvant ET alone (b) might benefit from the addition of chemotherapy, and high-risk HR+ patients who received chemoendocrine therapy (c) might benefit from the addition of CDK4/6 inhibitors. Finally, HR+ patients who received five years of adjuvant endocrine therapy (d) are candidates for extended endocrine treatment. A few assays have been shown to be associated with late recurrence and predict the benefit of extended endocrine therapy [22]. All plots in this figure are for the Providence cohort (the largest evaluation cohort).
@@
Traditional genomic assays are typically developed for specific patient subgroups, such as Oncotype DX for HR+ HER2− patients, and cannot be used in patients with less common molecular, histological, or clinical characteristics. Given that the AI test was developed using data from non-metastatic breast cancer patients without additional exclusion criteria, we can evaluate its prognostic accuracy in various subgroups. This includes patients with varying age and menopausal status, nodal status, tumor size, estrogen and HER2 receptor status, race, and administered of adjuvant therapy (Figure 3). Furthermore, our analysis indicates that the test score is consistent with known clinical factors that influence cancer prognosis, including higher scores in ER− and PR− patients, as well as patients with more advanced staging (both N and T) (Figure 5).
@@
NCCN guidelines also do not currently recommend any tests for the selection of adjuvant systemic therapy in most premenopausal HR+ patients (intermediate-risk node-negative and low- or intermediate-risk 1-3 LN+). The AI test reliably predicted DFI in HR+ premenopausal patients (if the menopausal status is unknown, we consider patients younger than 50 years as premenopausal), with a C-index of 0.655 [0.583-0.726] and a HR of 2.87 [1.81-4.55, p<0.01] (N= 677).
Furthermore, genomic assays have been reported to perform poorly in Black patients [24, 25]. The AI test reliably predicted DFI in Black patients (N= 299), with a C-index of 0.720 [0.65-0.79] and a HR 4.24 [3.27-5.50, p<0.01]. Among White patients (N= 2,461), the AI test achieved a C-index of 0.73 [0.69-0.79] and a HR of 3.44 [2.18-5.42, p<0.01].
Some prognostic genomic assays, including Oncotype DX, have also been shown to be predictive of adjuvant chemotherapy benefit [19, 26] in HR+ patients. We hypothesize that our test can identify patients who might benefit from adjuvant chemotherapy, CDK4/6 inhibitors, and extended endocrine therapy.
Figure 5: Relationship between predicted risk of recurrence and established prognostic factors. The top plot analyzes the Providence cohort with all invasive breast cancer subtypes to display the relationship with the ER/HER2 status. The bottom plot combines the Basel and Karmanos cohorts and illustrates the Oncotype score in relationship with the AI test score. Higher scores are observed in patients with HR− and HER2+ status and in patients with more advanced disease.
@@
Second, the AI test was prognostic in HR+ HER2− patients who received adjuvant chemoendocrine therapy (N=469, C-index: 0.70 [0.60-0.80], HR: 4.10 [2.09-8.02,p<0.01]). Patients who still have “residual” high risk while receiving both endocrine and chemotherapy are likely to benefit from the addition of CDK4/6 inhibitors.
Finally, the AI test was prognostic for late recurrence in HR+ HER2− patients (N=1158, C-index 0.79 [0.71-0.86], HR 8.44 [3.11-22.93, p<0.01]), suggesting that patients with high risk of cancer recurrence after 5 years might benefit from extended endocrine therapy.
To illustrate how our test can be used for clinical decision making in these subgroups, we defined high- and low-risk groups using the 80th percentile cutoff from our largest internal training dataset (it was giving us HR with significant p-value and similarly balanced high- and low-risk groups in all subgroups), verifying that high-risk patients have worse outcomes than low-risk patients. We apply this cutoff to the Providence cohort and compare outcomes between the low- and high-risk groups (Figure 4).
@@ 3 Discussion
In this study, we present a novel prognostic multi-modal AI test for invasive breast cancer. The overall performance of our test is robust across breast cancer patients. The test also outperformed Oncotype DX in three cohorts — Karmanos, Basel, and UChicago. Notably, the presented test is one of the few prognostic tools that have been clinically validated in triple-negative and HER2+ breast cancer patients. This capability, combined with strong performance regardless of nodal status and patient age, supports the viability of the AI test as a single tool to inform treatment decisions in all breast cancer patients.
The proposed AI test is designed to predict cancer recurrence risk, similar to genomic assays such as Oncotype DX and MammaPrint. However, just like these assays, it is not explicitly trained to model treatment effects. We expect that when validated using randomized prospective data, our prognostic test will show predictive capabilities in the same manner that the genomic assays do. That is, high-risk patients will benefit from more aggressive treatment options (such as adjuvant chemotherapy), while low-risk patients will have no benefit. While this study uses observational data and did not evaluate the test’s predictive capabilities (i.e., its ability to determine whether a patient is likely to benefit from a specific treatment), we present several potential clinical implications of using such a test in a clinical setting.
The AI test has the potential to refine clinical decision-making in several key areas. In HR+ patients who are candidates for adjuvant chemotherapy, the AI test was more accurate than Oncotype DX. This can potentially improve the selection of patients who can benefit from chemotherapy in addition to endocrine therapy. In a multivariate analysis, the AI test was shown to add independent prognostic information with respect to Oncotype and cancer grade, which indicates that it captures information that is strongly predictive and independent from existing clinical factors. Importantly, the AI test was prognostic in premenopausal HR+ patients, expanding patient groups who could be evaluated for cancer recurrence risk. Finally, we show that the AI test is prognostic in HR+ patients who received adjuvant chemoendocrine therapy, and identifies patients with high “residual” recurrence risk. We hypothesize that these patients might benefit from adjuvant CDK4/6 inhibitors. This use case is especially meaningful given recent FDA approval for ribociclib, which can be used in a wide group of HR+ patients, including node-negative cases. Identifying high-risk patients beyond basic clinical characteristics will be important in selecting patients who would benefit from CDK4/6 inhibitor treatment.
@@
The AI test was developed using Kestrel, an AI foundation model for digital pathology trained to extract morphological features from digitized pathology slides. Kestrel differs from the previous generation of feature extractors for digital pathology which relied on pre-specified “pathomic” features [32]. Kestrel, on the other hand, was trained with self-supervised learning, which does not require any pre-specification of features; rather, the self-supervised learning paradigm enables the model to learn the most salient features. While this sacrifices a priori explainability, the model can still be explained post hoc. Finally, because Kestrel was developed using a pan-cancer dataset, it can easily be applied to other clinical indications.
Applying self-supervised learning to develop prognostic/predictive models is an emerging research area. Spratt et al. demonstrated successful application of this approach to prostate cancer in a series of publications [33, 34]. The prostate test has been developed and validated using several randomized clinical trials. Garberis et al. developed an AI test for breast cancer based on morphological features extracted with contrastive learning which showed positive results [35]. However, that study was limited to standalone evaluation on a single ER+ HER2− patient cohort. Volinsky-Fremond et al. applied a self-supervised deep learning models to H&E-stained endometrial tumor slides, demonstrating performance superior to the current gold standard of combined pathological and molecular analysis, additionally finding their test predictive of chemotherapy benefit in the PORTEC-3 trial [36].
There are several analyses that we hope to carry out in the future to address potential limitations of this analysis. Most importantly, this study used retrospective observational data, which is not suitable for the evaluation of predictive capabilities. Also, the observational data may be biased given the small sample sizes of some cohorts. We mitigate this potential bias by analyzing several external datasets.
Integration of digital pathology-based tests into clinical workflow offers several advantages over genomic and other molecular prognostic and predictive tests. First, the turnaround time is significantly shorter, potentially allowing for more timely treatment decisions. Second, by utilizing standard H&E-stained slides and not requiring specialized wet lab tissue processing, the test is likely to be substantially cheaper than genomic assays, improving access to treatment selection testing in resource-limited settings. Third, our test can be performed on core biopsy slides, which allows for earlier risk assessment and might inform neoadjuvant therapy decisions. Finally, this digital approach preserves valuable tumor samples, which might be needed for deep sequencing in cases of disease progression.
@@ 4.1 Data
To execute this project we assembled a diverse collection of both public datasets (TCGA-BRCA [37, 38], METABRIC [39], BASIS [40]) and private datasets from several sources (NYU Langone Health, The Catholic University of Korea, Gundersen Health, National Center of Pathology in Lithuania, Breast Cancer Now Biobank, Cancer Research Malaysia, Omica.bio Cancer Atlas, Wales Cancer Biobank, Karmanos Cancer Institute, University Hospital Basel, and UChicago Medicine). All together we have 5162 slides from 4659 patients across 10 datasets for training, and 3632 slides from 3502 patients from 5 datasets for evaluation. This represents one of the largest data collection efforts for the purpose of predicting breast cancer recurrence. The breakdown of patient demographic, molecular, and clinical characteristics is described in Table 1.
Table 1: Datasets used for training and evaluation. All values are N (%) unless indicated otherwise.
@@ 4.2 Task and evaluation
Our primary task was to accurately predict the risk of cancer-related events. To accomplish this, we trained time-to-event models using clinical variables and digital pathology slides. For both modalities, we trained several models and ensembled their predictions, as detailed in Section 4.3.1. We then created a multi-modal model by averaging predictions of the two ensembles. We trained and evaluated the models using the definitions of endpoints in Table 3.
@@ 4.2.1 Performance metrics
Time-to-event models can be evaluated using a few different metrics that highlight various aspects of their performance. The two primary metrics we use are the concordance index (C-index) and hazard ratio (HR). Additionally, as we used several different datasets for evaluation, we aggregated results using random effects models. We refer to such aggregated metrics as pooled. In the following paragraphs, we justify the use of these metrics in our application. Detailed explanations of how the metrics are computed are in the Appendix.
@@
To evaluate the prognostic ability of our model, we used the concordance index (C-index). The C-index measures how well the predicted risk ranking of patients aligns with the actual order in which they experienced events. A C-index of 1.0 indicates that patients are perfectly ordered according to predicted risk, whereas a C-index of 0.5 suggests that the model’s ordering is no better than random. The advantage of the C-index is that it is easy to interpret, as it analogous to AUROC in how it is computed, while accommodating censoring. A detailed explanation of how the C-index is computed is in Appendix A.1.
Hazard ratio for a continuous score.
For a continuous score, the hazard ratio represents the relative increase or decrease in the risk of the event associated with a one-unit increase in that score. In other words, it quantifies how changes in the score influence the relative risk of the event happening over time. In our analysis, we compare a 0.2 unit increase in our score (which ranges from 0 to 1) to a 20 unit increase in the Oncotype score (which ranges from 0 to 100). A detailed explanation of how the hazard ratio is computed is in Appendix A.2.
Hazard ratio for a dichotomized score.
For a dichotomized score, the patients are stratified into two categories: high-risk and low-risk. The hazard ratio assesses the relative risk of experiencing an event in the two categories. For example, a HR of 2 indicates that the high-risk group is twice as likely to experience the event compared to the low-risk group. Hazard ratios come with associated p-values, computed with the Wald test, that evaluate whether the estimated hazard ratio for each covariate is statistically significantly different from 1.
@@
To assess the simultaneous impact of multiple variables on the outcome, we fit a multivariate Cox proportional hazards model with all the variables of interest. The Cox regression coefficients estimate the strength of association between each variable and the outcome, adjusted for other covariates. A coefficient greater than zero, or a hazard ratio exceeding 1 (as determined by the Wald test), suggests a positive association between the variable and the outcome after accounting for the influence of other factors. This analysis allows us to estimate the hazard ratio associated with different features after adjusting for the varying patient outcome distributions of different datasets. This analysis is performed to validate the association between our test and recurrence, even after adjusting for established prognostic factors such as cancer grade and Oncotype DX score. This analysis is used to assess whether our pathology model provides prognostic information independent of the clinical model.
@@ 4.2.2 Hyperparameter selection
In all instances, we selected the hyperparameters of the models through random search [41]. Unless otherwise explicitly stated we maximized the C-index. As we were tuning relatively few hyperparameters with moderately wide intervals, random search explores the hyperparameter space sufficiently well.
For pathology data, the performance of the models sampled in the hyperparameter search were estimated using modified multiple-source cross-validation [42]. That is, we had two sets of datasets $\mathcal{D}^{T}=\{D_{1},D_{2},\ldots,D_{J}\}$ and $\mathcal{D}^{V}=\{D_{J+1},D_{J+2},\ldots,D_{J+K}\}$. Datasets in $\mathcal{D}^{T}$ were reserved only for training, while $\mathcal{D}^{V}$ could be used for training or validation. For a given hyperparameter setting, we trained $K$ models, $m_{1}$, $m_{2}$, …, $m_{K}$. The training set for the $k^{\text{th}}$ model was $\mathcal{D}^{T}_{k}=\mathcal{D}^{T}\cup\mathcal{D}^{V}\backslash D_{J+k}$, and the validation set was $D_{J+k}\in\mathcal{D}^{V}$. The performances on the validation sets were averaged, weighted by the number of comparisons $c_{k}$ used in the computation of the validation C-index using the dataset $D_{J+k}$. This procedure was used for performance estimation and hyperparameter selection as formalized in Algorithm 1.
@@
In practice, to limit the variance in hyperparameter selection, for each set of hyperparameters, we trained and evaluated several models with different random seeds and averaged the validation performances.
In our implementation, when training the pathology models, we set $\mathcal{D}^{T}=\{\mathrm{Korea},\mathrm{Gundersen},\mathrm{Lithuania},\\ \mathrm{BCN},\mathrm{Malaysia},\mathrm{Karmanos},\mathrm{Basel},\mathrm{METABRIC},\mathrm{BASIS}\}$ and $\mathcal{D}^{V}=\{\mathrm{NYU},\mathrm{Omica},\mathrm{Wales}\}$.
As clinical models use a much lower dimensional input than pathology models and exhibit less variance, we adopted a simpler cross-validation procedure. We divided the data into $k$ splits, that is, $\mathcal{D}^{C}=\{D_{1}^{C},D_{2}^{C},\ldots,D_{K}^{C}\}$. For each iteration, we trained the models using $\mathcal{D}^{C}\backslash D_{k}^{C}$ and selected the hyperparameters that performs the best for this split based on validation performance on $D_{k}^{C}$. This procedure was used for performance estimation and hyperparameter selection as formalized in Algorithm 2.
@@
In our implementation, when training the clinical models, we trained using three splits $\mathcal{D}^{C}=\{D_{1}^{C},D_{2}^{C},D_{3}^{C}\}$. We set $D_{1}^{C}=\mathrm{NYU},D_{2}^{C}=\mathrm{METABRIC}$, and $D_{3}^{C}=\bigcup\{\mathrm{Korea},\mathrm{Gundersen},\mathrm{Lithuania},\mathrm{BCN},\mathrm{Malaysia},\\ \mathrm{Karmanos},\mathrm{Basel},\mathrm{BASIS},\mathrm{Wales}\}$.
4.3 The AI model
Our model learns and makes predictions using both clinical variables and digital pathology images. Internally, it is composed of many models whose predictions are ensembled in two rounds. First, the models are ensembled within their data modality. Then, the ensemble of clinical models is combined with the ensemble of digital pathology models. The intuition behind ensembling lies in the idea that different models make errors that are not perfectly correlated. Ensembling several models cancels out some of these errors, adding accuracy and robustness to predictions. An additional advantage of ensembling at the level of modalities is that for every test example, it is easy to interpret the contribution of the clinical data and the digital pathology data to the final prediction.

### v2

@@ 2.5 The AI test performs well in various clinically meaningful subgroups: potential clinical utility
Figure 4: The AI test is prognostic in clinically meaningful subgroups. All plots in this figure are for the Providence cohort (the largest evaluation cohort). a, In triple-negative breast cancer, there is an ongoing discussion about potential strategies to de-escalate the KEYNOTE-522 regimen, for example with shorter or anthracycline-free chemotherapy regimens or complete ommitance of adjuvant treatment. While some biomarkers, such as tumor-infiltrating lymphocytes, are associated with long-term outcomes and pathological complete response, there is still a need for a more robust assessment of systemic therapy benefits. b-d, In hormone receptor-positive patients, questions about treatment selection involve the addition of adjuvant chemotherapy, extended endocrine therapy (ET), and new agents, such as CDK4/6 inhibitors. We hypothesize that high-risk HR+ patients who received adjuvant ET alone (b) might benefit from the addition of chemotherapy, and high-risk HR+ patients who received chemoendocrine therapy (c) might benefit from the addition of CDK4/6 inhibitors. Finally, HR+ patients who received five years of adjuvant endocrine therapy (d) are candidates for extended endocrine treatment. A few assays have been shown to be associated with late recurrence and predict the benefit of extended endocrine therapy [27].
@@
Traditional genomic assays are typically developed for specific patient subgroups, such as Oncotype DX for HR+ HER2− patients, and cannot be used in patients with less common molecular, histological, or clinical characteristics. Given that the AI test was developed using data from non-metastatic breast cancer patients without additional exclusion criteria, we can evaluate its prognostic accuracy in various subgroups. This includes patients with varying age and menopausal status, nodal status, tumor size, estrogen and HER2 receptor status, race, and administered adjuvant therapy (Figure 3). Furthermore, our analysis indicates that the AI test score is associated with known clinical factors that influence cancer prognosis, including higher scores in ER− and PR− patients, as well as patients with more advanced staging (both N and T) (Figure 5).
@@
NCCN guidelines also do not currently recommend any tests for the selection of adjuvant systemic therapy in most premenopausal HR+ patients (intermediate-risk node-negative and low- or intermediate-risk 1-3 LN+). The AI test reliably predicted DFI in HR+ premenopausal patients (if the menopausal status is unknown, we consider patients younger than 50 years as premenopausal), with a C-index of 0.655 [0.583-0.726] and a HR of 2.87 [1.81-4.55, p=0.003] (n=677).
Furthermore, genomic assays have been reported to perform poorly in Black patients [29, 30]. As shown in Figure 3, the AI test reliably predicted DFI in Black patients (n=299), with a C-index of 0.720 [0.65-0.79] and a HR 4.24 [3.27-5.50, p=0.002]. Among White patients (n=2,461), the AI test achieved a C-index of 0.73 [0.69-0.79] and a HR of 3.44 [2.18-5.42, p=0.003].
Some prognostic genomic assays, including Oncotype DX, have also been shown to be predictive of adjuvant chemotherapy benefit [24, 31] in HR+ patients. We hypothesize that our test can, analogously, identify patients who might benefit from adjuvant chemotherapy, CDK4/6 inhibitors, and extended endocrine therapy.
Figure 5: Relationship between predicted risk of recurrence and established prognostic factors. The top plot analyzes the Providence cohort to display the relationship between the AI test score and the established factors such as ER/HER2 status and staging. The bottom plot combines the Basel, Karmanos and Chicago cohorts and illustrates the Oncotype score in relationship with the AI test score. Higher scores are observed in patients with HR− and HER2+ status and in patients with more advanced T or N staging. Higher AI test scores do not appear to be strongly associated with higher Oncotype DX risk, as indicated by the near-zero correlation ($R^{2}=0.02$) between our AI test scores and the Oncotype DX scores.
@@
Second, the AI test was prognostic in HR+ HER2− patients who received adjuvant chemoendocrine therapy (n=469, C-index: 0.70 [0.60-0.80], HR: 4.10 [2.09-8.02,p=0.007]). Patients who still have “residual” high risk while receiving both endocrine and chemotherapy are likely to benefit from the addition of CDK4/6 inhibitors.
Finally, the AI test was prognostic for late recurrence in HR+ HER2− patients (n=1158, C-index 0.79 [0.71-0.86], HR 8.45 [3.11-22.93, p<0.001]), suggesting that patients with higher risk of cancer recurrence after 5 years might benefit from extended endocrine therapy.
To illustrate the potential utility of our test in clinical decision-making, we defined high- and low-risk groups using the 80th percentile cutoff from our largest internal training dataset (NYU). For the analyses regarding treatment within HR+ HER2- subgroup, we used the 80th percentile within the HR+ HER2- subpopulation, while for the triple-negative breast cancer (TNBC) subgroup, we also applied the 80th percentile cutoff from its corresponding population. We selected this threshold as it was yielding statistically significant hazard ratios in our training set. Additionally, this threshold balanced the distribution of high- and low-risk groups across all subgroups, consistently demonstrating worse outcomes for high-risk patients compared to low-risk patients. As demonstrated in Figure 4, when we apply this cutoff to the Providence cohort, we see a significant difference in outcomes between the low- and high-risk groups.
@@ 3 Discussion
In this study, we present a novel prognostic multi-modal AI test for invasive breast cancer. The overall performance of our test is robust across breast cancer patients. We also compared the AI test to Oncotype DX in three cohorts where the Oncotype score is available. The test outperformed Oncotype DX in two cohorts, Karmanos and Basel, and performed as well as Oncotype DX in one additional cohort, UChicago. Notably, the presented test is one of the few prognostic tools that have been clinically validated in triple-negative and HER2+ breast cancer patients. This capability, combined with strong performance regardless of nodal status and patient age, supports the viability of the AI test as a single tool to inform treatment decisions in all breast cancer patients.
The proposed AI test is designed to predict cancer recurrence risk, similar to genomic assays such as Oncotype DX and MammaPrint. However, just like these assays, it is not explicitly trained to model treatment effects. We hypothesize that when validated using randomized prospective data, our prognostic test will show predictive capabilities in the same manner that the genomic assays do. That is, high-risk patients will benefit from more aggressive treatment options (such as adjuvant chemotherapy), while low-risk patients will have no benefit. While this study uses observational data and did not evaluate the test’s predictive capabilities (i.e., its ability to determine whether a patient is likely to benefit from a specific treatment), we present several potential clinical implications of using such a test in a clinical setting.
The AI test has the potential to refine clinical decision-making in several key areas. In HR+ patients who are candidates for adjuvant chemotherapy, the AI test was more accurate than Oncotype DX. This can potentially improve the selection of patients who can benefit from chemotherapy in addition to endocrine therapy. In a multivariate analysis, the AI test was shown to add independent prognostic information with respect to Oncotype and cancer grade, which indicates that it captures information that is strongly predictive and independent from these factors. Importantly, the AI test was prognostic in premenopausal HR+ patients, expanding patient groups who could be evaluated for cancer recurrence risk. Finally, we show that the AI test is prognostic in HR+ patients who received adjuvant chemoendocrine therapy, and identifies patients with high “residual” recurrence risk. We hypothesize that these patients might benefit from adjuvant CDK4/6 inhibitors. This use case is especially meaningful given recent FDA approval for ribociclib, which can be used in a wide group of HR+ patients, including node-negative cases. Identifying high-risk patients beyond basic clinical characteristics will be important in selecting patients who would benefit from CDK4/6 inhibitor treatment.
@@
Our test stands apart from existing genomic tests as it leverages pathology data and employs self-supervised learning. Previous generation of feature extractors for digital pathology relied on pre-specified “pathomic” features [37]. Self-supervised learning paradigm does not require any pre-specification of features; rather, it enables the model to learn the most salient features. Spratt et al. successfully applied this technique to prostate cancer, with their model validated through several randomized clinical trials [38, 39]. Similarly, Garberis et al. developed an AI test for breast cancer using contrastive learning, yielding strong results [40], though their study was limited to a single ER+ HER2− cohort. Volinsky-Fremond et al. extended the use of self-supervised deep learning to endometrial tumor slides, outperforming conventional combined pathological and molecular analyses and demonstrating predictive utility for chemotherapy benefit in the PORTEC-3 trial [41]. Our test is developed using Kestrel, an AI foundation model for digital pathology trained using self-supervised learning to extract morphological features from digitized pathology slides. As Kestrel was developed using a pan-cancer dataset, it showed promising results for a wide range of tasks across different clinical indications [17].
Integration of digital pathology-based tests into clinical workflow offers several advantages over genomic and other molecular prognostic and predictive tests. First, the turnaround time is significantly shorter, potentially allowing for more timely treatment decisions. Second, by utilizing standard H&E-stained slides and not requiring specialized wet lab tissue processing, the test is likely to be substantially cheaper than genomic assays, improving access to treatment selection testing and allowing for potentially pro-bono use in resource-limited settings. Third, our test can be performed on core biopsy slides, which allows for earlier risk assessment and might inform neoadjuvant therapy decisions. Finally, this digital approach preserves valuable tumor samples, which might be needed for deep sequencing in cases of disease progression.
In summary, this study underscores the potential of a digital pathology-based AI test in breast cancer prognosis. By delivering accurate, accessible, and cost-effective risk stratification in diverse patient populations, the test bridges significant gaps in current clinical workflows. With more rigorous validation through clinical trials, it could become an essential tool in personalized cancer care.
@@ 4.1 Data
The breakdown of patient demographic, molecular, and clinical characteristics of the training and test datasets is described in Table 1.
Table 1: Datasets used for training and evaluation. All values are N (%) unless indicated otherwise. Follow-up time statistics are computed taking into account all patients (censored and not censored).
@@ 4.2 Task and evaluation
Our primary task was to accurately predict the risk of cancer-related events. To accomplish this, we trained time-to-event models using clinical variables and digital pathology slides. For both modalities, we trained several models and ensembled their predictions, as detailed in Section 4.3.1. We then created a multi-modal model by averaging predictions of the two ensembles. We trained and evaluated the models using the endpoint definitions in Table A.3.
@@ 4.2.1 Performance metrics
The two primary metrics we use are the concordance index (C-index) and hazard ratio (HR). Apart from computing these metrics on each individual dataset, we use a random effects model to obtain pooled metrics which summarize performance over multiple datasets, as explained in Appendix A.3. In the following paragraphs, we justify the use of these metrics in our application.
@@
The C-index measures how well the predicted risk ranking of patients aligns with the actual order in which they experienced events or are censored. A C-index of 1.0 indicates that patients are perfectly ordered according to predicted risk, whereas a C-index of 0.5 indicates that the model’s ordering is no better than random. The advantage of the C-index is that it is easy to interpret, as it is analogous to AUROC in how it is computed, while accommodating censoring. A detailed explanation of how the C-index is computed is in Appendix A.1.
Hazard ratio of a continuous score.
For a continuous score, the hazard ratio represents the relative increase or decrease in the hazard of the event associated with a one-unit increase in that score. In our analysis, we compare a 0.2 unit increase in our AI test (which ranges from 0 to 1) to a 20 unit increase in the Oncotype score (which ranges from 0 to 100). A detailed explanation of how the hazard ratio is computed is in Appendix A.2.
Hazard ratio of a dichotomized score.
For a dichotomized score, the patients are stratified into two categories: high-risk and low-risk. The hazard ratio is computed using a Cox proportional hazard model with the dichotomized score as the only covariate, with the low-risk group as the reference. The p-values are computed using the Wald test.
@@
To assess the simultaneous impact of multiple variables on the outcome, we fit a multivariate Cox proportional hazards model with all variables of interest. For categorical variables in this analysis, we assign a binary indicator to all categories except one, which serves as reference.
We perform three types of multivariate analyses. In the first analysis, we include the AI test score alongside dataset indicators within the Cox model. This analysis allows us to estimate the hazard ratio associated with different features after adjusting for the varying patient outcome distributions in different datasets. We have a separate indicator variable for each dataset other than Basel, which is used as the reference group. In the second analysis, we have clinical and pathology scores within the AI test alongside the dataset indicator variables. This analysis is used to assess whether our pathology model provides prognostic information independent of the clinical model. In the final analysis, we use the AI test score along with Oncotype DX score, grade, race indicators and dataset indicators. For race, we have indicators for black, Asian and other, with white as the reference group. This analysis is performed to validate the association between our test and recurrence after adjusting for established prognostic factors.
@@ 4.2.2 Hyperparameter selection
In all instances, we selected the hyperparameters of the models through random search [42], using the training data. Unless otherwise explicitly stated we maximized the C-index.
In the hyperparameter search, the performance of pathology models was estimated using modified multiple-source cross-validation [43]. That is, we had two collections of datasets $\mathcal{D}^{T}=\{D_{1},D_{2},\ldots,D_{J}\}$ and $\mathcal{D}^{V}=\{D_{J+1},D_{J+2},\ldots,D_{J+K}\}$, where each $D$ in $\mathcal{D}^{T}$ and in $\mathcal{D}^{V}$ denotes a separate cohort. Datasets in $\mathcal{D}^{T}$ were reserved only for training, while $\mathcal{D}^{V}$ could be used for training or validation. For a given hyperparameter setting, we trained $K$ models, $m_{1}$, $m_{2}$, …, $m_{K}$. The training set for the $k^{\text{th}}$ model was $\mathcal{D}^{T}_{k}=\mathcal{D}^{T}\cup\mathcal{D}^{V}\backslash D_{J+k}$, and the validation set was $D_{J+k}\in\mathcal{D}^{V}$. The performances on the validation sets were averaged, weighted by the number of comparisons $c_{k}$ used in the calculation of the validation C-index using the dataset $D_{J+k}$. This procedure was used for performance estimation and hyperparameter selection as formalized in Algorithm 1. To limit the variance in hyperparameter selection, for each set of hyperparameters, we trained and evaluated several models with different random seeds and averaged the validation performances.
@@
In our implementation, when training the pathology models, we set $\mathcal{D}^{T}=\{\mathrm{Korea},\mathrm{Gundersen},\mathrm{Lithuania},\\ \mathrm{BCN},\mathrm{Malaysia},\mathrm{METABRIC},\mathrm{BASIS}\}$ and $\mathcal{D}^{V}=\{\mathrm{NYU},\mathrm{Omica},\mathrm{Wales}\}$.
As clinical models use much lower dimensional input than pathology models and their performance exhibits less variance, we adopted a simpler cross-validation procedure. We divided the data into $k$ splits, that is, $\mathcal{D}^{C}=\{D_{1}^{C},D_{2}^{C},\ldots,D_{K}^{C}\}$. For each iteration, we trained the models using $\mathcal{D}^{C}\backslash D_{k}^{C}$ and selected the hyperparameters that perform the best for this split based on validation performance on $D_{k}^{C}$. This procedure was used for performance evaluation and hyperparameter selection as formalized in Algorithm 2.
@@
In our implementation, when training the clinical models, we trained using three splits $\mathcal{D}^{C}=\{D_{1}^{C},D_{2}^{C},D_{3}^{C}\}$. We set $D_{1}^{C}=\mathrm{NYU},D_{2}^{C}=\mathrm{METABRIC}$, and $D_{3}^{C}=\bigcup\{\mathrm{Korea},\mathrm{Gundersen},\mathrm{Lithuania},\mathrm{BCN},\mathrm{Malaysia},\\ \mathrm{BASIS},\mathrm{Wales}\}$.
4.3 Models used to build the AI test
Our AI test makes predictions using both clinical variables and digital pathology images. Internally, it is composed of many models whose predictions are ensembled in two rounds. First, the models are ensembled within their data modality. Then, the ensemble of clinical models is combined with the ensemble of digital pathology models. The intuition behind ensembling lies in the idea that different models make errors that are not perfectly correlated. Ensembling several models cancels out some of these errors, adding accuracy and robustness to predictions. An additional advantage of ensembling at the level of modalities is that for every test example, it is easy to interpret the contribution of the clinical data and the digital pathology data to the final prediction.